# Estimates for trilinear and quadrilinear character sums

### v3

@@ 1.1. Set-up
We would like to emphasise that weights in $\mathfrak{Q}(\bm{\alpha},\bm{\beta})$ depend on variables from different products, which makes treatments of such sums much more difficult, as the standard smoothing technique does not immediately apply.
@@ 2.1. Bounds of multilinear character sums
for each positive integer $r$, where
@@
for each positive integer $r$, where
@@
for each positive integer $r$, where
@@
$\qquad+(KN)^{-\frac{1}{2}}p^{\frac{1}{2r}}+(MN)^{-\frac{1}{2}}+(LM)^{-\frac{1}{2}}+p^{-\frac{1}{2}}$
@@
(2.4) $\Delta_{3}=\widetilde{\Delta}_{3}+(KN)^{-\frac{1}{2}}p^{\frac{1}{2r}}+(MN)^{-\frac{1}{2}}+(LM)^{-\frac{1}{2}}+p^{-\frac{1}{2}}$
@@ 2.2. Consequences and applications
Fix $P$ a non-zero rational function in ${\mathbb{F}}_{p}(X)$ and $Q$ a non-zero rational function in ${\mathbb{F}}_{p}(X,Y)$. For every fixed $\eta>0$ there exists some $\kappa>0$, such that
@@
Another application of Theorem 2.7 concerns sums with the divisor function (see (1.8) below)
@@ 3.2. Bounds of some GCD sums
$G\leqslant\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}e_{1}f_{1}\ll A\\ d_{2}e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant KW+NU\end{subarray}}d_{1}d_{2}e_{1}e_{2}f_{1}f_{2}\mathop{\sum_{k\leqslant K}\sum_{n\leqslant N}\sum_{u\sim U}\sum_{w\sim W}}_{\operatorname{lcm}[f_{1},f_{2}]\mid(kw-nu)\neq 0}\mathop{\mathop{\sum\sum\sum\sum}_{\ell_{1},\ell_{2}\leqslant L,\leavevmode\nobreak\ m_{1},m_{2}\leqslant M}}_{\begin{subarray}{c}ak\ell_{1}+bm_{1}n\equiv 0\bmod{d_{1}e_{1}f_{1}}\\ a\ell_{2}u+bm_{2}w\equiv 0\bmod{d_{2}e_{2}f_{2}}\\ \ell_{\nu}\equiv 0\bmod{d_{\nu}},\,m_{\nu}\equiv 0\bmod{e_{\nu}},\ \nu=1,2\\ |ak\ell_{1}+bm_{1}n|\sim A,\leavevmode\nobreak\ |a\ell_{2}u+bm_{2}w|\sim B\end{subarray}}1.$
@@
$G\leqslant\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}e_{1}f_{1}\ll A\\ d_{2}e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant KW+NU\end{subarray}}d_{1}d_{2}e_{1}e_{2}f_{1}f_{2}\mathop{\sum_{k\leqslant K}\sum_{k\leqslant N}\sum_{u\sim U}\sum_{w\sim W}}_{\operatorname{lcm}[f_{1},f_{2}]\mid(kw-nu)\neq 0}\mathop{\mathop{\sum\sum\sum\sum}_{d_{1}\ell_{1},\,d_{2}\ell_{2}\leqslant L,\leavevmode\nobreak\ e_{1}m_{1},\,e_{2}m_{2}\leqslant M}}_{\begin{subarray}{c}ad_{1}\ell_{1}k+be_{1}m_{1}n\equiv 0\bmod{d_{1}e_{1}f_{1}}\\ ad_{2}\ell_{2}u+be_{2}m_{2}w\equiv 0\bmod{d_{2}e_{2}f_{2}}\\ |ad_{1}\ell_{1}k+be_{1}m_{1}n|\sim A,\leavevmode\nobreak\ |ad_{2}\ell_{2}u+be_{2}m_{2}w|\sim B\end{subarray}}1.$
@@
$G$ $\leqslant\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}e_{1}f_{1}\ll A\\ d_{2}e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant KW+NU\end{subarray}}d_{1}d_{2}e_{1}e_{2}f_{1}f_{2}\mathop{\sum_{k\leqslant K}\sum_{n\leqslant N}\sum_{u\sim U}\sum_{w\sim W}}_{\operatorname{lcm}[f_{1},f_{2}]\mid(kw-nu)\neq 0}\mathop{\mathop{\sum\sum\sum\sum}_{d_{1}\ell_{1},d_{2}\ell_{2}\leqslant L,\leavevmode\nobreak\ e_{1}m_{1},e_{2}m_{2}\leqslant M}}_{\begin{subarray}{c}e_{1}\mid ad_{1}\ell_{1}k,\leavevmode\nobreak\ bm_{1}n\equiv-ad_{1}\ell_{1}k/e_{1}\bmod{d_{1}f_{1}}\\ d_{2}\mid be_{2}m_{2}w,\leavevmode\nobreak\ a\ell_{2}u\equiv-be_{2}m_{2}w/d_{2}\bmod{e_{2}f_{2}}\\ |ad_{1}\ell_{1}k+be_{1}m_{1}n|\sim A,\leavevmode\nobreak\ |ad_{2}\ell_{2}u+be_{2}m_{2}w|\sim B\end{subarray}}1$
$\leqslant\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}f_{1}\ll A\\ e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant KW+NU\end{subarray}}d_{1}e_{2}f_{1}f_{2}\sup_{\delta\in{\mathbb{Z}}}\mathop{\sum_{k\leqslant K}\sum_{w\sim W}}_{kw\equiv\delta\bmod{\operatorname{lcm}[f_{1},f_{2}]}}\mathop{\sum\sum}_{\begin{subarray}{c}d_{2}\ll B/(e_{2}f_{2})\\ e_{1}\ll A/(d_{1}f_{1})\end{subarray}}d_{2}e_{1}$
@@
$G\ll AB\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}f_{1}\ll A\\ e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant KW+NU\end{subarray}}\sup_{\delta\in{\mathbb{Z}}}\mathop{\sum_{k\leqslant K}\sum_{w\sim W}}_{kw\equiv\delta\bmod{\operatorname{lcm}[f_{1},f_{2}]}}\mathop{\sum\sum}_{\begin{subarray}{c}d_{2}\ll B/(e_{2}f_{2})\\ e_{1}\ll A/(d_{1}f_{1})\end{subarray}}\mathop{\sum\sum}_{\begin{subarray}{c}\ell_{1}\leqslant L/d_{1},\leavevmode\nobreak\ m_{2}\leqslant M/e_{2}\\ d_{2}\mid be_{2}m_{2}w,\leavevmode\nobreak\ e_{1}\mid ad_{1}\ell_{1}k\end{subarray}}1.$
@@
$G\leqslant ABLM\Upsilon^{o(1)}\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}f_{1}\ll A\\ e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant KW+NU\end{subarray}}\frac{1}{d_{1}e_{2}}\sup_{\delta\in{\mathbb{Z}}}\mathop{\sum_{k\leqslant K}\sum_{w\sim W}}_{kw\equiv\delta\bmod{\operatorname{lcm}[f_{1},f_{2}]}}1.$
@@
$G$ $\leqslant ABLM)\Upsilon^{o(1)}\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}f_{1}\ll A\\ e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant KW+NU\end{subarray}}\frac{1}{d_{1}e_{2}}\left(\frac{KW}{\operatorname{lcm}[f_{1},f_{2}]}+1\right)$
@@ 4.2. Amplification
Note that we may further restrict our summation to triples $(k,m,n)$ with $p\nmid(ak+bmn)$ up to an error term at most
@@ 4.3. Regrouping, counting and Weil’s bound
${\mathcal{I}}_{1}\ll KM^{2}Np^{-\frac{1}{2r}}\qquad\mbox{and}\qquad{\mathcal{I}}_{2}\leqslant KM^{2}p^{-\frac{1}{2r}}(K+N)(1+\mathscr{L}_{1}p^{-1-\frac{1}{2r}})p^{o(1)}$
@@ 5.1. Cauchy–Schwarz inequality and amplification
with any fixed real-valued smooth function $W\in{\mathcal{C}}_{c}^{\infty}([-2,2])$, which majorates the characteristic function of the unit interval $[0,1]$.
@@ 5.2. Regrouping, counting and Weil’s bound
${\mathcal{J}}_{1}\ll K(MN)^{2}p^{-\frac{1}{r}}\qquad\mbox{and}\qquad{\mathcal{J}}_{2}\leqslant KN^{2}p^{-\frac{1}{r}}(K+M^{2})(1+\mathscr{L}p^{-1-\frac{1}{r}})^{2}p^{o(1)},$
@@
$T_{1}\leqslant\frac{1}{N}(KM^{2}N^{2})^{1-\frac{1}{2r}}(1+KM^{-2})^{\frac{1}{2r}}$ $\left(1+\mathscr{L}_{1}p^{-1-\frac{1}{r}}\right)^{\frac{1}{r}}$
@@ 6.1. Preliminary transformations
Assume $\|\bm{\alpha}\|_{\infty}\leqslant 1$. By the Cauchy–Schwarz inequality, we have
@@
for any smooth function $\varPhi$, which dominates the characteristic function of $[-1,1]$ and supported only inside of the interval $[-2,2]$. Squaring out and switching the order of summation, we get
(6.2) $Q\leqslant Q_{1}+KLMN(LM/p+1)p^{o(1)}$
@@
and, by (6.1), we obtain a non-trivial bound of $\mathfrak{Q}(\bm{\alpha},\bm{\beta})$ as soon as we improve (6.3). For all integers $(u,v,w)$, using that ${\mathbb{Z}}$ is invariant under shifts by integers, we can write
$Q_{1}$ $=\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}\ell_{1},\ell_{2}\leqslant L,\leavevmode\nobreak\ m_{1},m_{2}\leqslant M\\ \ell_{1}m_{2}\not\equiv\ell_{2}m_{1}\bmod p\end{subarray}}\mathop{\sum\sum}_{k,n\in{\mathbb{Z}}}\varPhi\left(\frac{k+wv}{K}\right)\varPhi\left(\frac{n+uv}{N}\right)\alpha_{\ell_{1},m_{1}}\overline{\alpha}_{\ell_{2},m_{2}}$
@@
(6.4) $UV=\frac{1}{4}K\qquad\mbox{and}\qquad WV=\frac{1}{4}N,$
@@
(6.5) $\begin{split}Q_{1}&\ll\frac{1}{UVW}\sum_{|k|\leqslant 2K}\sum_{|n|\leqslant 2N}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}\ell_{1},\ell_{2}\leqslant L,\leavevmode\nobreak\ m_{1},m_{2}\leqslant M\\ \ell_{1}m_{2}\not\equiv\ell_{2}m_{1}\bmod p\end{subarray}}\sum_{u\sim U}\sum_{w\sim W}\iint_{{\mathbb{R}}^{2}}|\widehat{\varPhi}(\xi)\widehat{\varPhi}(\eta)|\\ &\qquad\qquad\quad\times\biggl{|}\sum_{v\sim V}\mathrm{e}\left(\frac{uv\xi}{N}+\frac{wv\eta}{K}\right)\chi\left(\overline{(a\ell_{1}u+bm_{1}w)}(ak\ell_{1}+bm_{1}n)+v\right)\\ &\qquad\qquad\qquad\qquad\quad\times\overline{\chi}\left(\overline{(a\ell_{2}u+bm_{2}w)}(ak\ell_{2}+bm_{2}n)+v\right)\biggr{|}\mathrm{d}\xi\mathrm{d}\eta+\mathsf{Err}_{1},\end{split}$
@@
(6.7) $\mathsf{Err}_{1}\ll\frac{KL^{2}MNU}{UVW}\left(1+\frac{MW}{p}\right)p^{o(1)}\ll KL^{2}M\left(1+\frac{MN}{pV}\right)p^{o(1)}.$
@@
×$\qquad\qquad\qquad\times\sum_{u\sim U}\sum_{w\sim W}\biggl{|}\sum_{v\sim V}\mathrm{e}\left(\frac{v\xi}{N}+\frac{v\eta}{K}\right)\chi\left(\overline{(a\ell_{1}u+bm_{1}w)}(ak\ell_{1}+bm_{1}n)+v\right)$
@@
for some $\xi\in{\mathbb{R}}$, where $\mathsf{Err}_{1}$ satisfies (6.7). We now pull out $d=\gcd(ak,bn)$, for which we have $d\leqslant D:=\min\{|a|K,|b|N\}$. Hence, changing the variables $k\mapsto dk$ and $n\mapsto dn$, we obtain the inequality
(6.8) $Q_{1}\ll\frac{1}{UVW}\sum_{d\leqslant D}Q_{1}(d)\,+\mathsf{Err}_{1}$

### v4

@@ 1.1. Set-up
We would like to emphasize that weights in $\mathfrak{Q}(\bm{\alpha},\bm{\beta})$ depend on variables from different products, which makes treatments of such sums much more difficult, as the standard smoothing technique does not immediately apply.
@@ 2.1. Bounds of multilinear character sums
for each positive integer $r$, provided that $M>4p^{\frac{1}{2r}}$, where
@@
for each positive integer $r\geqslant 2$, provided that $N>4p^{\frac{1}{r}}$, where
@@
for each positive integer $r\geqslant 2$, provided that $K,N>4p^{\frac{1}{r}}$, where
@@
$\qquad+(KN)^{-\frac{1}{2}}p^{\frac{1}{2r}}+(MN)^{-\frac{1}{2}}p^{\frac{1}{2r}}+(LM)^{-\frac{1}{2}}+p^{-\frac{1}{2}}$
@@
(2.4) $\Delta_{3}=\widetilde{\Delta}_{3}+(KN)^{-\frac{1}{2}}p^{\frac{1}{2r}}+(MN)^{-\frac{1}{2}}p^{\frac{1}{2r}}+(LM)^{-\frac{1}{2}}+p^{-\frac{1}{2}}$
@@ 2.2. Consequences and applications
Fix a rational function $P$ in ${\mathbb{F}}_{p}(X)$ and a rational function $Q$ in ${\mathbb{F}}_{p}(X,Y)$. For every fixed $\eta>0$ there exists some $\kappa>0$, such that
@@
Another application of Theorem 2.7 concerns sums with the divisor function (see (1.8) above)
@@ 3.2. Bounds of some GCD sums
$G\leqslant\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}e_{1}f_{1}\ll A\\ d_{2}e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant 2(KW+NU)\end{subarray}}d_{1}d_{2}e_{1}e_{2}f_{1}f_{2}\mathop{\sum_{k\leqslant K}\sum_{n\leqslant N}\sum_{u\sim U}\sum_{w\sim W}}_{\operatorname{lcm}[f_{1},f_{2}]\mid(kw-nu)\neq 0}\mathop{\mathop{\sum\sum\sum\sum}_{\ell_{1},\ell_{2}\leqslant L,\leavevmode\nobreak\ m_{1},m_{2}\leqslant M}}_{\begin{subarray}{c}ak\ell_{1}+bm_{1}n\equiv 0\bmod{d_{1}e_{1}f_{1}}\\ a\ell_{2}u+bm_{2}w\equiv 0\bmod{d_{2}e_{2}f_{2}}\\ \ell_{\nu}\equiv 0\bmod{d_{\nu}},\,m_{\nu}\equiv 0\bmod{e_{\nu}},\ \nu=1,2\\ |ak\ell_{1}+bm_{1}n|\sim A,\leavevmode\nobreak\ |a\ell_{2}u+bm_{2}w|\sim B\end{subarray}}1.$
@@
$G\leqslant\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}e_{1}f_{1}\ll A\\ d_{2}e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant 2(KW+NU)\end{subarray}}d_{1}d_{2}e_{1}e_{2}f_{1}f_{2}\mathop{\sum_{k\leqslant K}\sum_{k\leqslant N}\sum_{u\sim U}\sum_{w\sim W}}_{\operatorname{lcm}[f_{1},f_{2}]\mid(kw-nu)\neq 0}\mathop{\mathop{\sum\sum\sum\sum}_{d_{1}\ell_{1},\,d_{2}\ell_{2}\leqslant L,\leavevmode\nobreak\ e_{1}m_{1},\,e_{2}m_{2}\leqslant M}}_{\begin{subarray}{c}ad_{1}\ell_{1}k+be_{1}m_{1}n\equiv 0\bmod{d_{1}e_{1}f_{1}}\\ ad_{2}\ell_{2}u+be_{2}m_{2}w\equiv 0\bmod{d_{2}e_{2}f_{2}}\\ |ad_{1}\ell_{1}k+be_{1}m_{1}n|\sim A,\leavevmode\nobreak\ |ad_{2}\ell_{2}u+be_{2}m_{2}w|\sim B\end{subarray}}1.$
@@
$G$ $\leqslant\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}e_{1}f_{1}\ll A\\ d_{2}e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant 2(KW+NU)\end{subarray}}d_{1}d_{2}e_{1}e_{2}f_{1}f_{2}\mathop{\sum_{k\leqslant K}\sum_{n\leqslant N}\sum_{u\sim U}\sum_{w\sim W}}_{\operatorname{lcm}[f_{1},f_{2}]\mid(kw-nu)\neq 0}\mathop{\mathop{\sum\sum\sum\sum}_{d_{1}\ell_{1},d_{2}\ell_{2}\leqslant L,\leavevmode\nobreak\ e_{1}m_{1},e_{2}m_{2}\leqslant M}}_{\begin{subarray}{c}e_{1}\mid ad_{1}\ell_{1}k,\leavevmode\nobreak\ bm_{1}n\equiv-ad_{1}\ell_{1}k/e_{1}\bmod{d_{1}f_{1}}\\ d_{2}\mid be_{2}m_{2}w,\leavevmode\nobreak\ a\ell_{2}u\equiv-be_{2}m_{2}w/d_{2}\bmod{e_{2}f_{2}}\\ |ad_{1}\ell_{1}k+be_{1}m_{1}n|\sim A,\leavevmode\nobreak\ |ad_{2}\ell_{2}u+be_{2}m_{2}w|\sim B\end{subarray}}1$
$\leqslant\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}f_{1}\ll A\\ e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant 2(KW+NU)\end{subarray}}d_{1}e_{2}f_{1}f_{2}\sup_{\delta\in{\mathbb{Z}}}\mathop{\sum_{k\leqslant K}\sum_{w\sim W}}_{kw\equiv\delta\bmod{\operatorname{lcm}[f_{1},f_{2}]}}\mathop{\sum\sum}_{\begin{subarray}{c}d_{2}\ll B/(e_{2}f_{2})\\ e_{1}\ll A/(d_{1}f_{1})\end{subarray}}d_{2}e_{1}$
@@
$G\ll AB\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}f_{1}\ll A\\ e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant 2(KW+NU)\end{subarray}}\sup_{\delta\in{\mathbb{Z}}}\mathop{\sum_{k\leqslant K}\sum_{w\sim W}}_{kw\equiv\delta\bmod{\operatorname{lcm}[f_{1},f_{2}]}}\mathop{\sum\sum}_{\begin{subarray}{c}d_{2}\ll B/(e_{2}f_{2})\\ e_{1}\ll A/(d_{1}f_{1})\end{subarray}}\mathop{\sum\sum}_{\begin{subarray}{c}\ell_{1}\leqslant L/d_{1},\leavevmode\nobreak\ m_{2}\leqslant M/e_{2}\\ d_{2}\mid be_{2}m_{2}w,\leavevmode\nobreak\ e_{1}\mid ad_{1}\ell_{1}k\end{subarray}}1.$
@@
$G\leqslant ABLM\Upsilon^{o(1)}\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}f_{1}\ll A\\ e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant 2(KW+NU)\end{subarray}}\frac{1}{d_{1}e_{2}}\sup_{\delta\in{\mathbb{Z}}}\mathop{\sum_{k\leqslant K}\sum_{w\sim W}}_{kw\equiv\delta\bmod{\operatorname{lcm}[f_{1},f_{2}]}}1.$
@@
$G$ $\leqslant ABLM\Upsilon^{o(1)}\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}f_{1}\ll A\\ e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant 2(KW+NU)\end{subarray}}\frac{1}{d_{1}e_{2}}\left(\frac{KW}{\operatorname{lcm}[f_{1},f_{2}]}+1\right)$
@@ 4.2. Amplification
We may further restrict our summation to triples $(k,m,n)$ with $p\nmid(ak+bmn)$ up to an error term at most
@@ 4.3. Regrouping, counting and Weil’s bound
${\mathcal{I}}_{1}\ll KM^{2}Np^{-\frac{1}{2r}}\qquad\mbox{and}\qquad{\mathcal{I}}_{2}\leqslant KM^{2}p^{-\frac{1}{2r}}(K+N)(1+\mathscr{L}_{1}MNp^{-1-\frac{1}{2r}})p^{o(1)}$
@@ 5.1. Cauchy–Schwarz inequality and amplification
with any fixed real-valued smooth function $W\in{\mathcal{C}}_{c}^{\infty}([-2,2])$, which majorizes the characteristic function of the unit interval $[0,1]$.
@@ 5.2. Regrouping, counting and Weil’s bound
${\mathcal{J}}_{1}\ll K(MN)^{2}p^{-\frac{1}{r}}\qquad\mbox{and}\qquad{\mathcal{J}}_{2}\leqslant KN^{2}p^{-\frac{1}{r}}(K+M^{2})(1+\mathscr{L}_{1}MNp^{-1-\frac{1}{r}})^{2}p^{o(1)},$
@@
$T_{1}\leqslant\frac{1}{N}(KM^{2}N^{2})^{1-\frac{1}{2r}}(1+KM^{-2})^{\frac{1}{2r}}$ $(1+\mathscr{L}_{1}MNp^{-1-\frac{1}{r}})^{\frac{1}{r}}$
@@ 6.1. Preliminary transformations
Assume $\|\bm{\alpha}\|_{\infty}\leqslant 1$ and $\gcd(a,b)=1$ without loss of generality. By periodicity, we also assume that $21\leqslant|a|,|b|<p/$ 2. By the Cauchy–Schwarz inequality, we have
@@
for any smooth function $\varPhi$, which dominates the characteristic function of $[-1,1]$ and is supported only inside the interval $[-2,2]$. Squaring out and switching the order of summation, we get
(6.2) $Q\leqslant|Q_{1}|+KLMN(LM/p+1)p^{o(1)}$
@@
and, by (6.1), we obtain a non-trivial bound of $\mathfrak{Q}(\bm{\alpha},\bm{\beta})$ as soon as we improve (6.3). For all integers $u,v,w$, using that ${\mathbb{Z}}$ is invariant under shifts by integers, we can write
$Q_{1}$ $=\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}\ell_{1},\ell_{2}\leqslant L,\leavevmode\nobreak\ m_{1},m_{2}\leqslant M\\ \ell_{1}m_{2}\not\equiv\ell_{2}m_{1}\bmod p\end{subarray}}\mathop{\sum\sum}_{k,n\in{\mathbb{Z}}}\varPhi\left(\frac{k+uv}{K}\right)\varPhi\left(\frac{n+wv}{N}\right)\alpha_{\ell_{1},m_{1}}\overline{\alpha}_{\ell_{2},m_{2}}$
@@
(6.4) $U,V,W\geqslant 1,\ \ \ \ UV=\frac{1}{4}K\qquad\mbox{and}\qquad WV=\frac{1}{4}N,$
@@
(6.5) $\begin{split}Q_{1}&\ll\frac{1}{UVW}\sum_{|k|\leqslant 2K}\sum_{|n|\leqslant 2N}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}\ell_{1},\ell_{2}\leqslant L,\leavevmode\nobreak\ m_{1},m_{2}\leqslant M\\ \ell_{1}m_{2}\not\equiv\ell_{2}m_{1}\bmod p\end{subarray}}\sum_{u\sim U}\sum_{w\sim W}\iint_{{\mathbb{R}}^{2}}|\widehat{\varPhi}(\xi)\widehat{\varPhi}(\eta)|\\ &\qquad\qquad\quad\times\biggl{|}\sum_{v\sim V}\mathrm{e}\left(\frac{uv\xi}{K}+\frac{wv\eta}{N}\right)\chi\left(\overline{(a\ell_{1}u+bm_{1}w)}(ak\ell_{1}+bm_{1}n)+v\right)\\ &\qquad\qquad\qquad\qquad\quad\times\overline{\chi}\left(\overline{(a\ell_{2}u+bm_{2}w)}(ak\ell_{2}+bm_{2}n)+v\right)\biggr{|}\mathrm{d}\xi\mathrm{d}\eta+\mathsf{Err}_{1},\end{split}$
@@
(6.7) $\mathsf{Err}_{1}\leqslant\frac{KL^{2}MNUV}{UVW}\left(1+\frac{MW}{p}\right)p^{o(1)}\leqslant KL^{2}M\left(V+\frac{MN}{p}\right)p^{o(1)}.$
@@
×$\qquad\qquad\qquad\times\sum_{u\sim U}\sum_{w\sim W}\biggl{|}\sum_{v\sim V}\mathrm{e}\left(\frac{v\xi}{K}+\frac{v\eta}{N}\right)\chi\left(\overline{(a\ell_{1}u+bm_{1}w)}(ak\ell_{1}+bm_{1}n)+v\right)$
@@
for some $\xi\in{\mathbb{R}}$, where $\mathsf{Err}_{1}$ satisfies (6.7). We now pull out the gcd of $ak$ and $bn$ for latter purpose. To this end, we put
$d=\gcd(ak,bn),\leavevmode\nobreak\ \leavevmode\nobreak\ d_{1}=\frac{d}{\gcd(a,d)},\leavevmode\nobreak\ \leavevmode\nobreak\ d_{2}=\frac{d}{\gcd(b,d)},\leavevmode\nobreak\ \leavevmode\nobreak\ a^{*}=\frac{a}{\gcd(a,d)},\leavevmode\nobreak\ \leavevmode\nobreak\ b^{*}=\frac{b}{\gcd(b,d)}.$
We observe that $p\nmid d$, $\gcd(a,d)=\gcd(a,n)$, $\gcd(b,d)=\gcd(b,k)$, and $d\leqslant D$ with
$D=\min\{|a|K,|b|N\}\leavevmode\nobreak\ \leavevmode\nobreak\ (<p^{2}).$
Note that $d\mid ak$ implies $d_{1}\mid k$, and similarly, $d\mid bn$ implies $d_{2}\mid n$. Hence, we have the relations $d_{1}\mid k$, $d_{2}\mid n$, $ad_{1}=a^{*}d$ and $bd_{2}=b^{*}d$. Therefore, changing the variables
(6.8) $k\mapsto d_{1}k,\ \ \ n\mapsto d_{2}n,$
we obtain the inequality
(6.9) $Q_{1}\ll\frac{1}{UVW}\sum_{\begin{subarray}{c}d\leqslant D\\ p\nmid d\end{subarray}}Q_{1}(d)\,+\mathsf{Err}_{1}$